a) Suppose the own price elasticity of demand for good X is -2, its income elasticity of demand is 3, its advertising elasticity of demand is 4, and the cross-price elasticity of demand between it and good Y is -6. Determine how much the consumption of this good will change if:

i. The price of good X increases by 5 percent.
ii. The price of good Y increase by 10 percent.
iii. Advertising decreases by 2 percent.
iv. Income falls by 3 percent.
1
Expert's answer
2018-08-07T15:40:09-0400
i. The quantity demanded decreases by 2*5=10%
ii. The quantity demanded decreases by 6*10=60%
iii. The quantity demanded decreases by 2*4=8%
iv. The quantity demanded decreases by 3*3=9%